OpenStudy (anonymous):

Which of the following is equivalent to 73 = 343? log3343 = 7 log7343 = 3 log37 = 343 log73 = 343

OpenStudy (anonymous):

I mean 7^3 = 343

OpenStudy (anonymous):

there must be some mistake here because i am fairly sure that \[73\neq 343\]

OpenStudy (anonymous):

ooooooh ok \[b^x=y\iff \log_b(y)=x\]

OpenStudy (anonymous):

you need to be able so switch back and forth more or less automatically \[2^3=8\iff \log_2(8)=3\] \[7^2=49\iff \log_7(49)=2\] \[10^6=1,000,000\iff \log_{10}(1,000,000)=6\] \[

OpenStudy (anonymous):

\[\log_{7} 7^{3}=3\] 7^3=343

OpenStudy (anonymous):

Thanks Again

OpenStudy (anonymous):

ok